# Hierarchical Budget Policy Optimization for Adaptive Reasoning

### v1

@@ 1 Introduction
Advances in large reasoning models have led to impressive performance on complex reasoning tasks through chain-of-thought methodologies (Lea, ; Guo et al., ). However, these models exhibit fundamental inefficiency: they generate unnecessarily long reasoning chains even for simple problems, sometimes consuming thousands of tokens for basic arithmetic (Chen et al., b; a). This phenomenon reveals a fundamental misalignment, as current reasoning models lack the ability to adapt their computational effort to the actual complexity of problems.
@@
Recent empirical findings challenge the conventional belief that longer reasoning always leads to better outcomes. Research shows that models can maintain competitive accuracy even without intermediate steps (Ma et al., ), and in some cases, shorter reasoning paths perform comparably or even better on simpler tasks (Li et al., ). This is further supported by stark variations in optimal reasoning lengths across tasks. For instance, L1 (Aggarwal & Welleck, ) achieves peak performance with  1,100 tokens on GSM8K, but requires over 3,000 tokens on OlympiadBench. Such heterogeneity highlights a key insight: the computational requirements for effective reasoning are inherently problem-dependent, yet current models apply uniform reasoning strategies regardless of task complexity.
@@
To address these inefficiencies, an increasing number of studies aim to improve the inference efficiency of reasoning models. Current approaches fall into two primary categories. Length-controlled methods directly constrain generation through explicit mechanisms: prompts like ”think for n tokens” and corresponding length-control rewards in L1 (Aggarwal & Welleck, ); progressively limits on the model’s reasoning space during training in ThinkPrune (Hou et al., ); and Scalable Chain of Thoughts (Xu et al., ) enforces budget constraints through forced termination. Reward-based approaches incorporate length penalties into training objectives: HAPO (Huang et al., ) leverages history-aware optimization to track minimal sufficient reasoning lengths, while ”Think When You Need” (Jiang et al., ) employs pairwise comparison rewards to trade off between quality and brevity. While effective at reducing token usage, these methods share a key limitation: they prioritize efficiency at the cost of accuracy, lacking mechanisms for models to decide autonomously when longer or shorter reasoning is appropriate.
@@
We identify two key challenges that hinder existing methods from achieving genuine reasoning efficiency. First, length penalties introduce systematic training biases that impair reasoning capabilities. In standard reinforcement learning settings (Guo et al., ), correct solutions receive equal rewards regardless of length, allowing for unbiased exploration. However, length penalties disrupt this balance by consistently favoring shorter outputs, leading policies to gradually abandon long-reasoning strategies (Hou et al., ; Huang et al., ; Lou et al., ). Second, static efficiency constraints fail to capture the continuous nature of reasoning complexity. Even adaptive methods rely on coarse mechanisms, such as binary think/no-think decisions (Zhang et al., a; Fang et al., ) or fixed confidence thresholds (Qiao et al., ), which overlook the nuanced relationship between problem characteristics and computational requirements.
@@
These limitations raise a fundamental question: rather than enforcing uniform constraints, can models learn differentiated reasoning strategies through structured exploration? This question motivates our study of hierarchical budget exploration, where efficiency emerges not from rigid length control but from guiding models through structured exploration within budget-constrained subspaces.
@@
We propose Hierarchical Budget Policy Optimization (HBPO), illustrated in the bottom-left part of Figure 1, a reinforcement learning framework that enables models to learn problem-specific reasoning strategies while retaining their ability to perform complex reasoning. The core idea is to partition the exploration space into multiple budget-constrained subgroups, allowing models to preserve reasoning diversity and uncover natural alignments between problem characteristics and required computational effort. Specifically, HBPO employs a hierarchical sampling strategy that partitions rollout samples into subgroups, each governed by a distinct token budgets. We implement this by inserting length prompts (e.g., ”I will answer the question within n tokens”) after the reasoning tag, thereby constructing multiple exploration spaces with budgets ranging from 512 to 2560 tokens. Unlike uniform sampling, this structure encourages the model to explore both concise and extended reasoning paths throughout training, effectively mitigating the systematic degradation of reasoning capabilities caused by global length penalties.
@@
  • We introduce Hierarchical Budget Policy Optimization (HBPO), a reinforcement learning framework that partitions the exploration space into budget-constrained hierarchies with differentiated rewards, preserving reasoning diversity while enabling adaptive resource allocation.
@@ 2.1 Efficient Reasoning
Recent advances in reasoning models have spurred various efforts to reduce computational overhead while preserving performance. Existing approaches can be broadly categorized into three types: Length-constrained methods explicitly restrict generation through predefined mechanisms. For example, L1 (Aggarwal & Welleck, ) introduces token budget prompts with corresponding rewards; ThinkPrune (Hou et al., ) progressively tightens constraints via iterative training; and Scalable Chain of Thoughts (Xu et al., ) separates the thinking and solution phases, each with its budget. While effective in limiting token usage, these methods require manual budget specification and lack adaptability to varying problem complexity. Reward-based methods incorporate efficiency into training objectives more implicitly. HAPO (Huang et al., ) incentivizes concise reasoning by tracking minimal correct response lengths, while “Think When You Need” (Jiang et al., ) balances brevity and quality through pairwise comparisons and adaptive target lengths. These approaches offer finer control but still impose global objectives across diverse problem types, limiting flexibility. Training-free approaches (Qiao et al., ; Muennighoff et al., ; Yang et al., ) intervene at inference time through symbolic control tokens or confidence-based early stopping. While cost-effective, these methods are heuristic-driven and lack learning-based adaptation. Despite their differences, all these approaches share a fundamental limitation: they treat efficiency as a uniform constraint, overlooking the fact that optimal reasoning length varies significantly with problem complexity.
@@ 2.2 Adaptive Reasoning
Recognizing heterogeneous reasoning requirements, recent work explores adaptive strategies that adjust computational effort based on problem characteristics. Binary mode selection represents the most common approach, with models choosing between thinking and non-thinking modes (Lou et al., ; Zhang et al., a; Fang et al., ). These methods employ various techniques including selective loss masking, simplified mode definitions, and decoupled optimization to prevent mode collapse. Multi-stage training strategies (Jiang et al., ; Tu et al., ; Zhang et al., b) use sophisticated reward designs and batch-level balancing to achieve better mode distributions. Beyond binary selection, multi-modal approaches define richer reasoning taxonomies: ARM (Wu et al., ) uses four modes with adaptive scaling, while PATS (Wang et al., ) enables step-level switching between complexity levels. Some methods introduce auxiliary components like regression models for mode prediction (Liang et al., ) or self-budgeting mechanisms (Li et al., ). While these adaptive approaches demonstrate significant efficiency gains, they operate within discrete categories rather than enabling continuous adaptation. Complex multi-stage procedures and predefined mode taxonomies limit their flexibility and generalization. In contrast, our hierarchical budget exploration framework enables continuous adaptation through a unified policy optimization process. Without relying on manually defined modes or external modules, our approach allows the model to learn problem-specific reasoning depths, leading to emergent adaptive behavior that naturally aligns computational effort with problem complexity.
@@ 3 Method
We present Hierarchical Budget Policy Optimization (HBPO), as shown in Figure 2, which extends the Group Relative Policy Optimization (GRPO) (Guo et al., ) framework to enable adaptive reasoning through structured exploration. The core innovation lies in partitioning the exploration space into budget-constrained hierarchies and designing differentiated reward mechanisms that preserve reasoning diversity. We first introduce the hierarchical rollout strategy (Section 3.1), then detail the budget-aware reward design (Section 3.2)), and finally describe the training procedure (Section 3.3)).
@@ 4.1 Experimental Setup
We evaluate HBPO on mathematical reasoning tasks using the DeepScaleR dataset (Luo et al., 2025) for training, which comprises 40K high-quality mathematical problems from AIME, AMC, Omni-Math (Gao et al., 2024), and STILL (Min et al., ). We employ two base models: DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI, 2025) and DeepScaleR-Preview-1.5B (Luo et al., 2025).
@@
We implement HBPO using the VeRL framework (Sheng et al., 2024) with a context window of 4,096 tokens during training. Following DAPO (Yu et al., ), we set clipping thresholds $\epsilon_{\text{high}}=0.28$ and $\epsilon_{\text{low}}=0.2$, with KL divergence disabled to encourage exploration. Training proceeds for one epoch (629 steps) with a learning rate of $10^{-6}$ and batch size of 64. For hierarchical exploration, we generate 16 rollouts per query, partitioned equally into 4 subgroups with budget constraints $\mathcal{B}={512,1024,2048,2560}$ tokens.
@@
We evaluate on four mathematical reasoning benchmarks of increasing difficulty: GSM8K (Cobbe et al., 2021), Math500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), and AIME25. Following standard practice (Guo et al., ), we use temperature $T=0.6$, $top\_p=0.95$, and maximum context length of 32,768 tokens. We report pass@1 accuracy and average token usage under two evaluation settings: (1) natural reasoning where models freely determine their computational effort, and (2) efficiency prompting using “I will answer the question with minimal tokens” after <think> to guide models toward efficient responses.
@@
We compare against several state-of-the-art efficient reasoning methods: (1) TLMREArora & Zanette , which adds length penalties to the RL objective with hyperparameter $\alpha$ controlling the penalty strength; (2) AdaptThinkZhang et al. (a) and AutoThinkTu et al. , which enable binary think/no-think mode selection; (3) L1-MaxAggarwal & Welleck , which uses two-stage RL with explicit length targets. These methods represent different approaches to reasoning efficiency: global penalties (TLMRE), discrete mode selection (AdaptThink, AutoThink), and explicit control (L1-Max).
@@ 6 Conclusion
We introduced Hierarchical Budget Policy Optimization (HBPO), a framework that enables reasoning models to achieve efficient computation without sacrificing capability. By maintaining diverse exploration through budget-constrained hierarchies and budget-aware rewards, HBPO prevents the exploration collapse and an optimized allocation of the length budget. Our experiments demonstrate that models trained with HBPO significantly reduce inference costs while improving performance, exhibiting adaptive behavior that naturally matches computational effort to problem complexity.

### v2

@@ 1 Introduction
Advances in large reasoning models have led to impressive performance on complex reasoning tasks through chain-of-thought methodologies (OpenAI, 2024; DeepSeek-AI, 2025b). However, these models exhibit fundamental inefficiency: they generate unnecessarily long reasoning chains even for simple problems, sometimes consuming thousands of tokens for basic arithmetic (Chen et al., 2025; 2024). This phenomenon reveals a fundamental misalignment, as current reasoning models lack the ability to adapt their computational effort to the actual complexity of problems.
@@
Recent empirical findings challenge the conventional belief that longer reasoning always leads to better outcomes. Research shows that models can maintain competitive accuracy even without intermediate steps (Ma et al., 2025), and in some cases, shorter reasoning paths perform comparably or even better on simpler tasks (Li et al., 2025). This is further supported by stark variations in optimal reasoning lengths across tasks. For instance, L1 (Aggarwal & Welleck, 2025) achieves peak performance with  1,100 tokens on GSM8K, but requires over 3,000 tokens on OlympiadBench. Such heterogeneity highlights a key insight: the computational requirements for effective reasoning are inherently problem-dependent, yet current models apply uniform reasoning strategies regardless of task complexity.
@@
To address these inefficiencies, an increasing number of studies aim to improve the inference efficiency of reasoning models. Current approaches fall into two primary categories. Length-controlled methods directly constrain generation through explicit mechanisms: prompts like “think for n tokens” and corresponding length-control rewards in L1 (Aggarwal & Welleck, 2025); progressively limits on the model’s reasoning space during training in ThinkPrune (Hou et al., 2025); and Scalable Chain of Thoughts (Xu et al., 2025) enforces budget constraints through forced termination. Reward-based approaches incorporate length penalties into training objectives: HAPO (Huang et al., 2025) leverages history-aware optimization to track minimal sufficient reasoning lengths, while “Think When You Need” (Jiang et al., 2025) employs pairwise comparison rewards to trade off between quality and brevity. While effective at reducing token usage, these methods share a key limitation: they prioritize efficiency at the cost of accuracy, lacking mechanisms for models to decide autonomously when longer or shorter reasoning is appropriate.
@@
We identify two key challenges that hinder existing methods from achieving genuine reasoning efficiency. First, length penalties introduce systematic training biases that impair reasoning capabilities. In standard reinforcement learning settings (DeepSeek-AI, 2025b), correct solutions receive equal rewards regardless of length, allowing for unbiased exploration. However, length penalties disrupt this balance by consistently favoring shorter outputs, leading policies to gradually abandon long-reasoning strategies (Hou et al., 2025; Huang et al., 2025; Lou et al., 2025). Second, static efficiency constraints fail to capture the continuous nature of reasoning complexity. Even adaptive methods rely on coarse mechanisms, such as binary think/no-think decisions (Zhang et al., 2025a; Fang et al., 2025) or fixed confidence thresholds (Qiao et al., 2025), which overlook the nuanced relationship between problem characteristics and computational requirements.
@@
These limitations raise a fundamental question: rather than enforcing uniform constraints, can models learn differentiated reasoning strategies through structured exploration? This question motivates our study of hierarchical budget exploration, where efficiency emerges not from rigid control but from structured exploration within budget-constrained subspaces.
@@
We propose Hierarchical Budget Policy Optimization (HBPO), illustrated in the bottom-left part of Figure 1, a reinforcement learning framework that enables models to learn problem-specific reasoning strategies while retaining their ability to perform complex reasoning. The core idea is to partition the exploration space into multiple budget-constrained subgroups, allowing models to preserve reasoning diversity and uncover natural alignments between problem characteristics and required computational effort. Specifically, HBPO employs a hierarchical sampling strategy that partitions rollout samples into subgroups, each governed by a distinct token budgets. We implement this by inserting length prompts (e.g., “I will answer the question within n tokens”) after the reasoning tag, thereby constructing multiple exploration spaces with budgets ranging from 512 to 2560 tokens. Unlike uniform sampling, this structure encourages the model to explore both concise and extended reasoning paths throughout training, effectively mitigating the systematic degradation of reasoning capabilities caused by global length penalties.
@@
  • We introduce Hierarchical Budget Policy Optimization, a reinforcement learning framework that partitions the exploration space into budget-constrained hierarchies with differentiated rewards, preserving reasoning diversity while enabling adaptive resource allocation.
@@ 2.1 Efficient Reasoning
Recent advances in reasoning models have spurred various efforts to reduce computational overhead while preserving performance. Existing approaches can be broadly categorized into three types: Length-constrained methods explicitly restrict generation through predefined mechanisms. For example, L1 (Aggarwal & Welleck, 2025) introduces token budget prompts with corresponding rewards; ThinkPrune (Hou et al., 2025) progressively tightens constraints via iterative training; and Scalable Chain of Thoughts (Xu et al., 2025) separates the thinking and solution phases, each with its budget. While effective in limiting token usage, these methods require manual budget specification and lack adaptability to varying problem complexity. Reward-based methods incorporate efficiency into training objectives more implicitly. HAPO (Huang et al., 2025) incentivizes concise reasoning by tracking minimal correct response lengths, while “Think When You Need” (Jiang et al., 2025) balances brevity and quality through pairwise comparisons and adaptive target lengths. These approaches offer finer control but still impose global objectives across diverse problem types, limiting flexibility. Training-free approaches (Muennighoff et al., 2025; Yang et al., 2025) intervene at inference time through symbolic control tokens or confidence-based early stopping. While cost-effective, these methods are heuristic-driven and lack learning-based adaptation. Despite their differences, all these approaches share a fundamental limitation: they treat efficiency as a uniform constraint, overlooking the fact that optimal reasoning length varies significantly with problem complexity.
@@ 2.2 Adaptive Reasoning
Recognizing heterogeneous reasoning requirements, recent work explores adaptive strategies that adjust computational effort based on problem characteristics. Binary mode selection represents the most common approach, with models choosing between thinking and non-thinking modes (Lou et al., 2025; Zhang et al., 2025a; Fang et al., 2025). These methods employ various techniques including selective loss masking, simplified mode definitions, and decoupled optimization to prevent mode collapse. Multi-stage training strategies (Jiang et al., 2025; Tu et al., 2025; Zhang et al., 2025b) use sophisticated reward designs and batch-level balancing to achieve better mode distributions. Beyond binary selection, multi-modal approaches define richer reasoning taxonomies: ARM (Wu et al., 2025) uses four modes with adaptive scaling, while PATS (Wang et al., 2025) enables step-level switching between complexity levels. Some methods introduce auxiliary components like regression models for mode prediction (Liang et al., 2025) or self-budgeting mechanisms (Li et al., 2025). While these adaptive approaches demonstrate significant efficiency gains, they operate within discrete categories rather than enabling continuous adaptation. Complex multi-stage procedures and predefined mode taxonomies limit their flexibility and generalization. In contrast, our hierarchical budget exploration framework enables continuous adaptation through a unified policy optimization process. Without relying on manually defined modes or external modules, our approach allows the model to learn problem-specific reasoning depths, leading to emergent adaptive behavior that naturally aligns computational effort with problem complexity.
@@ 3 Method
We present Hierarchical Budget Policy Optimization, as shown in Figure 2, which extends the Group Relative Policy Optimization (GRPO) (DeepSeek-AI, 2025b) framework to enable adaptive reasoning through structured exploration. The core innovation lies in partitioning the exploration space into budget-constrained hierarchies and designing differentiated reward mechanisms that preserve reasoning diversity. We first introduce the hierarchical rollout strategy (Section 3.1), then detail the budget-aware reward design (Section 3.2)), and finally describe the training procedure (Section 3.3)).
@@ 4.1 Experimental Setup
We evaluate HBPO on mathematical reasoning tasks using the DeepScaleR dataset (Luo et al., 2025) for training, which comprises 40K high-quality mathematical problems from AIME, AMC, Omni-Math (Gao et al., 2025), and STILL (Min et al., 2024). We employ two base models: DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI, 2025a) and DeepScaleR-Preview-1.5B (Luo et al., 2025).
@@
We implement HBPO using the VeRL framework (Sheng et al., 2024) with a context window of 4,096 tokens during training. Following DAPO (Yu et al., 2025), we set clipping thresholds $\epsilon_{\text{high}}=0.28$ and $\epsilon_{\text{low}}=0.2$, with KL divergence disabled to encourage exploration. Training proceeds for one epoch (629 steps) with a learning rate of $10^{-6}$ and batch size of 64. For hierarchical exploration, we generate 16 rollouts per query, partitioned equally into 4 subgroups with budget constraints $\mathcal{B}={512,1024,2048,2560}$ tokens.
@@
We evaluate on four mathematical reasoning benchmarks of increasing difficulty: GSM8K (Cobbe et al., 2021), Math500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), and AIME25. Following standard practice (DeepSeek-AI, 2025b), we use temperature $T=0.6$, $top\_p=0.95$, and maximum context length of 32,768 tokens. We report pass@1 accuracy and average token usage under two evaluation settings: (1) natural reasoning where models freely determine their computational effort, and (2) efficiency prompting using “I will answer the question with minimal tokens” after <think> to guide models toward efficient responses.
@@
We compare against several state-of-the-art efficient reasoning methods: (1) TLMRE(Arora & Zanette, 2025), which adds length penalties to the RL objective with hyperparameter $\alpha$ controlling the penalty strength; (2) AdaptThink(Zhang et al., 2025a) and AutoThink(Tu et al., 2025), which enable binary think/no-think mode selection; (3) L1-Max(Aggarwal & Welleck, 2025), which uses two-stage RL with explicit length targets. These methods represent different approaches to reasoning efficiency: global penalties (TLMRE), discrete mode selection (AdaptThink, AutoThink), and explicit control (L1-Max).
@@ 6 Conclusion
We introduced Hierarchical Budget Policy Optimization, a framework that enables reasoning models to achieve efficient computation without sacrificing capability. By maintaining diverse exploration through budget-constrained hierarchies and budget-aware rewards, HBPO prevents the exploration collapse and an optimized allocation of the length budget. Our experiments demonstrate that models trained with HBPO significantly reduce inference costs while improving performance, exhibiting adaptive behavior that naturally matches computational effort to problem complexity.